Shift report 20000824 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 24 2000 - 08:23:10 EDT

  • Next message: http account: "Shift report 20000824 08:00-16:00"

    ------------------------------------------------------------
    Date: Thu Aug 24 08:23:09 2000, Shift: 00:00-08:00
    Supervisor: ykl, with: b2
    Log book#5 pages 4 - 6
    
    Shift Summary:
    ------------------------------------------------------------
    1.Continuous beam was available except for an hour and half
    for a beam dump at 3:54.
    2. Run 2422 continuing from the previous shift gave 574K of
    Trigger 1 which accomplished 800K of Trigger 1 with D5 
    polarity A. So the D5 polarity is changed to B and the run
    2423 was started with 350A for D5.
    3. The anode and field of T1 tripped frequently (within
    a few minutes) even after the field HV was raised to 450.
    The server tsir crashed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 24 2000 - 08:26:23 EDT